Student number: Name: Lithium has two stable isotopes with natural abundances of 92.41%ai Li). How many 3. atoms of Li are there in a 1.54 g sample of lithium oxide? of ) and 7.59% A) 1.87 ×1020 B) 3.001020 C) 7.32 x102 D) 4.71 x10
Answer #1

average atomicmass = (7-Li mass*percentage + 6-Li mass*percentage)/100

7-Li mass = 7 u

7-Li percentage = 92.41%

6-Li mass = 6 u

6-Li percentage = 7.59%

average Li-atomicmass = ((7*92.41)+(6*7.59))/(92.41+7.59)

                      = 6.924 u

No of mol of Li = w/at.wt

                = 1.54/29.9

                = 0.0515 mol


1 Li2O = 2 li atoms

No of mol of Li atoms = 0.0515*2 = 0.103 mol

No of mol of 6-Li atoms = 0.103*7.59/100 = 0.00782 mol

No of atoms of 6-Li = 0.00782 *6.023*10^23

                     = 4.71*10^21 atoms

answer: D